# Deep Phase Coded Image Prior

### 05 Apr 2024 v1

@@ Abstract
Phase-coded imaging is a computational imaging method designed to tackle tasks such as passive depth estimation and extended depth of field (EDOF) using depth cues inserted during image capture. Most of the current deep learning-based methods for depth estimation or all-in-focus imaging require a training dataset with high-quality depth maps and an optimal focus point at infinity for all-in-focus images. Such datasets are difficult to create, usually synthetic, and require external graphic programs. We propose a new method named ”Deep Phase Coded Image Prior” (DPCIP) for jointly recovering the depth map and all-in-focus image from a coded-phase image using solely the captured image and the optical information of the imaging system. Our approach does not depend on any specific dataset and surpasses prior supervised techniques utilizing the same imaging system. This improvement is achieved through the utilization of a problem formulation based on implicit neural representation (INR) and deep image prior (DIP). Due to our zero-shot method, we overcome the barrier of acquiring accurate ground-truth data of depth maps and all-in-focus images for each new phase-coded system introduced. This allows focusing mainly on developing the imaging system, and not on ground-truth data collection.
@@ 1 Introduction
Phase coded imaging is a method in computational imaging, where a camera is equipped with an optical phase mask that provides unambiguous depth-related color characteristics for the captured image. While single-image passive depth estimation is an ill-posed problem, in phase-coded imaging, a phase mask is designed to create a different defocus blur for each color channel of the image. The depth cues, encoded by the phase mask, can later be decoded by a deep neural network for passive depth estimation or extended depth-of-field.
@@
Prior works decode depth cues via a supervised learning approach, where a synthetic dataset with accurate depth maps and all-in-focus images is used, and a neural network is trained to extract the depth map and/or all-in-focus image from the blurred (coded) image. A major drawback of these methods is the need for a large and accurate dataset for training. Capturing such a dataset can be quite complex and may necessitate the use of external sensors or graphic rendering software.
Another issue is generalization, namely, a deep learning model trained on synthetic data may not generalize well to real-world data. Moreover, even a model trained on real-world data might suffer from a domain shift between the training and test sets.
@@
To address the aforementioned limitations, we propose a new approach inspired by prior image restoration techniques utilizing Deep Image Prior[1]. We formulate the tasks of single-image passive depth estimation and extraction of all-in-focus images as a single inverse problem. An implicit generator with an input code is optimized to jointly create an all-in-focus image and a depth map. The loss function of the generator relies on a Differential Camera Model (DCM) that simulates phase-coded imaging from a given depth map and all-in-focus image. Thus, our loss term compares the input phase-coded image to the one generated by the DCM based on neural network reconstruction. Note that the learning process is end-to-end. Thus, the gradients propagate all the way back to the generator to improve the mapping between the input code and the reconstructed depth map and all-in-focus image such that the DCM output matches the input. At the end of the learning process, we discard the output and just keep the depth map and an all-in-focus RGB image. Figure 1 illustrates this process.
@@
Contribution. Our contributions may be summarized as (i) proposing a method to extract both depth estimation and all-in-focus images in a zero-shot manner using a known DCM, which reduces the need to capture a curated dataset for training; and (ii) outperforming prior works trained in a supervised fashion for the same imaging system.
@@
Figure 1: System overview. We follow a neural field flow, where an implicit generator first maps the encoded input image into reconstruction representations (in our case an all-in-focus image and a depth map). In the second part, A Differential Camera Model (DCM), a simulated differential forward (acquisition) process of a phased-coded imaging system, provides the acquired image. Both parameters in the generation part, and in the DCM part, are optimized based on the loss defined between the estimated phase-coded image and the input image.
@@ 2.1 Coded Aperture Imaging
Coded aperture imaging systems employ dedicated optical masks to encode light. Such imaging can decode and extract visual cues in order to reconstruct a passive depth estimation of the acquired image. Levin et al. [2] use an amplitude-coded mask such that objects at different depths exhibit a distinctive spectral structure. Haim et al. [3] proposed a ring phase mask to generate depth-dependent PSF filters for encoding the depth cues. They proposed a fully-convolutional neural network to learn the decoding process to generate the depth map. A follow-up work by Gil et al. [4] leverages the encoding process in a stereo vision setup where conventional cameras fail. Chang et al. [5] proposed a depth-dependent image formation where a neural network learns to extract a depth map from a simulated acquired image produced from an all-in-focus image, a binned depth map and the lens parameters. Wu et al. [6] proposed an end-to-end architecture consisting of two parts: An optical layer simulates depth-dependent PSFs given a learnable phase mask, and a reconstruction network, which estimates depth from the coded image.
@@ 2.2 Deep Image Prior
GT Input Mono from [4] DPCIP
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 2: A qualitative comparison of depth map estimation reveals that our method generates significantly more precise depth maps compared to the mono network from [4], which tends to excessively smooth the estimated maps, resulting in lower accuracy. Note that the depths are clipped to the physical range of the imaging system.
@@ 2.3 Implicit Neural Representation
Implicit neural representation is an area of research that studies the abilities of neural networks to map input coordinates to target values. This includes 2D images, 3D shapes, SDF maps and many more. It was demonstrated in [23] that representing the input coordinates as Fourier-features with a tuneable bandwidth enables a simple MLP to generate complex target domains such as images and 3D shapes while preserving their high-level details. In Neural Radiance Fields (NeRF) [24], implicit functions with Fourier features are used for synthesizing novel views of a 3D scene from sparse 2D images. SIREN [25] showed that by changing the activation function of a simple MLP from the frequently used ReLU to a periodic activation (for example a sine activation), the network can represent the spatial and temporal derivatives using a simple input grid to successfully recover a wide range of target domains (images, videos, 3D surfaces, etc.). Following NeRF, improvements were proposed [26, 27, 28] in terms of convergance speed compared to the original NeRF by utilizing a structure for the implicit model. Methods were also proposed [29, 30, 31] to tackle the scaling and aliasing problems that the original NeRF suffered from. Yu et al. [32] suggested a method to reduce the need for many calibrated images from sparse views as supervision for NeRF. Mildenhall et al. [33] suggested a method for rendering the raw images before the ISP in order to achieve much wider rendering abilities such as control on tone mapping. It was also proposed [34] that Gaussian activation can perform better than sine activation while having the property of being continuous everywhere and at every derivative. SAPE [35] and BACON [36] demonstrate how a combination of a coordinate network along with a limitation of its frequency spectrum can achieve a multi-scale representation of the target domain.
@@
TABLE I: A quantitative comparison of depth estimation, results are depth RMSE error [m] (normal-$\downarrow$). Our method outperforms the mono network from [4] in depth estimation by 2 orders of magnitude on the same imaging system, even though the Monocular depth estimation network from [4] was trained in a supervised manner on the same dataset, this is likely due to the fact that dataset is small and the model is prone to overfitting.
@@
3 Deep Phase Coded Image Prior
We consider an imaging system capable of extracting simultaneously an all-in-focus RGB image and a pixel-level depth map with a single image capture. Our goal is to achieve single image depth estimation and extended depth of field (EDOF) simultaneously relying solely on the captured image without any training data. We achieve this via end-to-end joint optimization of a generation part (a U-Net shaped neural network) and a forward process part, namely, a differentiable approximation of the acquisition process of the phase-coded imaging system - DCM in short.
As shown in Figure 1 We formulate our problem as an implicit neural representation problem where our system is composed of two major components: The first one is an implicit generator based on the model introduced in DIP [1]. It is an encoder-decoder model such that the encoder has 5 levels and each level has 2 convolutional layers with a depth of 128 and the downsampling between the levels is done by stride. The decoder has also 5 levels with 2 convolutional layers with a depth of 128 followed by bilinear upsampling. At each level the model has skip connections as a convolutional layer with a depth of 16. This creates a generator whose learnable parameter is the mapping between an input code and a pair consisting of an all-in-focus image and a pixel-level depth map. The second component is a differntiable simulation of the phase-coded optical imaging (DCM). Given a pair of an all-in-focus image and a depth map, it generates the acquired image from a phase-coded imaging system.
We simulate the coded-phase acquisition process as done in [3, 7], i.e., assuming the optical parameters are known (lens parameters and the focus point), we can recover the out of focus (OOF) condition (denoted as $\psi$) using the following equation:
where $R$ is the radius of the exit pupil (circular aperture is assumed), $\lambda$ is the illumination wavelength, $z_{img}$ is the sensor plane location for an object at a nominal position $z_{n}$, and $z_{i}$ is the ideal location for an image plane with respect to an object at location $z_{0}$.
From Eq. 1 we can derive that a focused object is located at $\psi=0$, and an out of focus object will be located at other values of $\psi$. $\psi$ ranges between $[-4,10]$. With the knowledge of the focus point and the recovered $\psi$ values for each pixel, we can recover the absolute depth for each pixel in the scene.
Our method and simulations are based on [3, 7], where we have a ring mask with pre-defined depth-dependent PSF kernels. Since the acquisition process is not differentiable, we used a differentiable approximation, in a coarse to fine manner. Instead of creating a mask that corresponds to a particular depth and convolving the specific depth with the associated pre-calculated PSF kernels, inspired by [9, 5], our DCM first creates several acquired images from a fixed set of depths ($\psi$) values and then linearly interpolates each pixel with the closest depth-dependent blur images.
The differentiable variant allows us to optimize the whole system in a self-supervised end-to-end manner, where the generator first maps the input encoding to a pair of an all-in-focus image and a pixel-level depth map that are fed as inputs to the DCM in order to reproduce the acquired image taken by the camera. The joint optimization of producing both an all-in-focus image and a pixel-level depth map forces an alignment between them and yields better results compared to previous works where a model was trained to produce just a depth map or just an all-in-focus image.
To optimize our neural network weights to map a coded input to pixel values we used the non differential acquisition simulation to produce coded-phase images to act as our supervision. Our objective function is to the minimize difference between a given image and the forward process output. Similarly to [13], we first use L2 as our reconstruction loss and switch to SSIM [37] after a fixed number of iterations. We can formulate our objective as follows:
$\theta^{*}=\operatornamewithlimits{argmin}_{x}\ell(h(f_{\theta}(z))-y),$ (2)
GT Input EDOF DPCIP

### 09 Mar 2025 v2

@@ Abstract
Phase-coded imaging is a computational imaging scheme designed to tackle tasks such as passive depth estimation and extended depth of field (EDOF) by inserting depth cues during image capture. Most of the current deep learning-based methods for depth estimation or all-in-focus imaging require a training dataset with high-quality depth maps and an optimal focus point at infinity for all-in-focus images. These datasets are difficult to create, often synthetic, and require external graphic software. Thus, limited datasets are being used, limiting the diversity of applicable domains. We propose a new method named “Deep Phase Coded Image Prior” (DPCIP) for jointly recovering the depth map and all-in-focus image from a coded-phase image using solely the captured image and the optical information of the imaging system. Our approach does not depend on any specific dataset and surpasses prior supervised techniques utilizing the same imaging system. This improvement is achieved through the utilization of a problem formulation based on implicit neural representation (INR) and deep image prior (DIP). Due to our zero-shot method, we overcome the barrier of acquiring accurate ground-truth data of depth maps and all-in-focus images for each new phase-coded system introduced. This allows focusing mainly on developing the imaging system, and not on ground-truth data collection.
@@ 1 Introduction
Phase-coded imaging is a method in computational imaging, where a camera is equipped with an optical phase mask that provides unambiguous depth-related color characteristics for the captured image. While single-image passive depth estimation is an ill-posed problem, in phase-coded imaging, a phase mask is designed to create a different defocus blur for each color channel of the image. The depth cues, encoded by the phase mask, can later be decoded by a deep neural network for passive depth estimation or extended depth-of-field.
@@
Prior works decode depth cues using a supervised learning approach, which involves training a neural network on a dataset with an accurate ground-truth to extract the depth map and/or all-in-focus image from the blurred (coded) image. . A major drawback of these methods is the need for a large and accurate dataset for training. Capturing such datasets can be quite complex and may require external sensors or graphic rendering software. While synthetic datasets using graphic software are easier to create than real-world datasets with external depth sensors, they require expertise, and their ability to generalize to real-world applications is uncertain. Real-world datasets are limited in quantity. Thus, there are two options: rely on known datasets, which limits the problem domain, or create a dedicated dataset, which is limited and may suffer from a domain gap between training and test sets.
@@
To address the aforementioned limitations, we propose a new approach inspired by prior image restoration techniques that utilize Deep Image Prior [1]. We formulate the tasks of single-image passive depth estimation and extraction of all-in-focus images as a unified inverse problem. An implicit generator is optimized to jointly create an all-in-focus image and a depth map from a given input code. The outputs from the generator are then passed through a Differential Camera Model (DCM) that simulates phase-coded imaging from a given depth map and all-in-focus image, resulting in a blurred image with depth cues. Our reconstruction loss compares the given phase-coded image to the one generated by the DCM. Since our camera model is differential the gradients propagate all the way back to the generator to improve the mapping between the input code and the depth map and all-in-focus image such that, after passing through the DCM, the produced output matches the input. At the end of the learning process, we discard the final output and retain the intermediate results: the depth map and an all-in-focus RGB image. Figure 1 illustrates this process.
@@
Contribution. Our contributions can be summarized as: (i) proposing a method to extract both depth estimation and all-in-focus images in a zero-shot manner using a known DCM, reducing the need for a curated dataset for training; and (ii) outperforming prior supervised works for the same imaging system. Our code is available at https://github.com/NimrodShabtay/DPCIP.
@@
Figure 1: System overview. We follow a neural field flow, where an implicit generator first maps the encoded input image into reconstruction representations (in our case an all-in-focus image and a depth map). In the second part, A Differential Camera Model (DCM) simulates the acquisition process of a phased-coded imaging system, provides the acquired image. The gradients from the reconstruction loss flow back to the generator forcing it to produce accurate intermediate outputs to match the acquired image after passing through the DCM. In the bottom row we illustrate the actual acquisition process compared to our neural field flow.
@@ 2.1 Coded Aperture Imaging
Coded aperture imaging systems employ dedicated optical masks to encode light. Such imaging can decode and extract visual cues in order to reconstruct a passive depth estimation of the acquired image. Levin et al. [2] use an amplitude-coded mask such that objects at different depths exhibit a distinctive spectral structure. Haim et al. [3] proposed a ring phase mask to generate depth-dependent PSF filters for encoding the depth cues. They proposed a fully-convolutional neural network to learn the decoding process to generate the depth map. A follow-up work by Gil et al. [4] leverages the encoding process in a stereo vision setup where conventional cameras fail. Chang et al. [5] proposed a depth-dependent image formation where a neural network learns to extract a depth map from a simulated acquired image produced from an all-in-focus image, a binned depth map and the lens parameters. Wu et al. [6] proposed an end-to-end architecture consisting of two parts: An optical layer that simulates depth-dependent PSFs given a learnable phase mask, and a reconstruction network, which estimates depth from the coded image.
@@ 2.3 Implicit Neural Representation
Implicit neural representation is an area of research that studies the abilities of neural networks to map input coordinates to target values. This includes 2D images, 3D shapes, SDF maps and many more. It was demonstrated in [23] that representing the input coordinates as Fourier-features with a tuneable bandwidth enables a simple MLP to generate complex target domains such as images and 3D shapes while preserving their high-level details. In Neural Radiance Fields (NeRF) [24], implicit functions with Fourier features are used for synthesizing novel views of a 3D scene from sparse 2D images. SIREN [25] showed that by changing the activation function of a simple MLP from the frequently used ReLU to a periodic activation (for example a sine activation), the network can represent the spatial and temporal derivatives using a simple input grid to successfully recover a wide range of target domains (images, videos, 3D surfaces, etc.). Following NeRF, improvements were proposed [26, 27, 28] in terms of convergence speed compared to the original NeRF by utilizing a structure for the implicit model. Methods were also proposed [29, 30, 31] to tackle the scaling and aliasing problems that the original NeRF suffered from. Yu et al. [32] suggested a method to reduce the need for many calibrated images from sparse views as supervision for NeRF. Mildenhall et al. [33] suggested a method for rendering the raw images before the ISP in order to achieve much wider rendering abilities such as control on tone mapping. It was also proposed [34] that Gaussian activation can perform better than sine activation while having the property of being continuous everywhere and at every derivative. SAPE [35] and BACON [36] demonstrate how a combination of a coordinate network along with a limitation of its frequency spectrum can achieve a multi-scale representation of the target domain.
3 Deep Phase Coded Image Prior
We consider an imaging system capable of extracting simultaneously an all-in-focus RGB image and a pixel-level depth map with a single image capture. Our goal is to achieve single image depth estimation and extended depth of field (EDOF) simultaneously relying solely on the captured image without any training data.Eliminating the need for a large dataset is a major advantage over supervised methods, which use synthetic or real-world datasets from limited domains and are sensitive to domain shift and generalization [37, 38, 39], thus they are sensitive to domain shift and generalization. We achieve this via a joint optimization of a generation part (a U-Net shaped neural network) and a forward process part, namely, a differentiable approximation of the acquisition process of the phase-coded imaging system - DCM in short.
As shown in Figure 1, we formulate our problem as an implicit neural representation problem where our system is composed of two major components: an implicit generator and a Differentiable Camera Model (DCM).
3.1 Phase Mask
We simulate the coded-phase acquisition process as done in [3, 7]. The proposed mask is a two-ring phase mask designed to manipulate the PSF of the imaging system based on color and depth. For a color RGB sensor, three separate behaviors can be generated, such that at each depth of the scene, a different channel (R/G/B) is in focus while the others are not. Figure 2 describes the hardware setup of our system.
Assuming the optical parameters are known (lens parameters and the focus point), we can recover the out of focus (OOF) parameter (denoted as $\psi$) from [40]. For the case of a circular exit pupil with radius $R$, the following equation defines the OOF parameter:
where $\lambda$ is the illumination wavelength, $z_{img}$ is the sensor plane location for an object at a nominal position $z_{n}$, and $z_{i}$ is the ideal location for an image plane with respect to an object at location $z_{0}$.
From Eq. 1 we can derive that a focused object is located at $\psi=0$, and an out of focus object will be located at other values of $\psi$. While defocus has a symmetry around the focus point, it is not possible to generate a symmetric phase mask for multiple colors ($\pi$ phase for one color is a different phase for another). Therefore, we enhance the color-dependent effects, and create an extremely nonsymmetric mask, which behaves very differently for each color. The asymmetric design can help us extend the depth of field (DOF) for each color separately (i.e. for every depth one color channel will provide good contrast, one medium contrast and one poor). These depth-color cues will serve also as the depth hints. Since closer objects will be larger in the final image, they can ‘suffer’ from more defocus so we take the longer part of the antisymmetric domain towards the closer objects. The range values [-4, 10] is a design choice, given the hardware and optics components.
Refer to caption Refer to caption Refer to caption
Figure 2: Hardware setup. Our setup is composed of a phase mask embedded inside the lens on the aperture plane. On the left, the optical components are placed seperately. In the middle, the phase mask is located on the aperture plane. On the right, you can see a schematic view of the phase mask pattern.
3.2 Implicit Generator
The implicit generator is based on the model introduced in DIP [1]. It is an encoder-decoder model such that the encoder has 5 levels and each level has 2 convolutional layers with a depth of 128 and 3x3 spatial kernels. The downsampling between the levels is done by stride. The decoder has also 5 levels with 2 convolutional layers with a depth of 128 and 3x3 spatial kernels followed by bilinear upsampling. At each level the model has skip connections as a 1x1 convolutional layer with a depth of 16. This creates a generator whose learnable parameter is the mapping between an input code and a pair consisting of an all-in-focus image and a pixel-level depth map. Like in DIP, we define the input code to be uniform random noise, $X_{input}\sim\mathcal{U}[0,1]$, where $X_{input}\in\mathbb{R}^{D\times H\times W}$.
3.3 Differentiable Camera Model (DCM)
The DCM is a differentiable simulation of the phase-coded optical imaging. Given a pair of an all-in-focus image and a depth map, it generates the acquired image from a phase-coded imaging system. Our method and simulations are based on [3, 7], where we have a ring mask with pre-defined depth-dependent PSF kernels. Since the acquisition process has several non-differentiable operations (i.e. ‘argmax’), we used a differentiable approximation, inspired by [9, 5] - The DCM takes the generated sharp RGB image and creates several blur images based on a fixed set of defocus ($/psi$) values. Each image is a single defocus blur corresponding to a particular depth. Then, with the generated depth map, the DCM linearly interpolates each pixel based on the set of blur images it created to produce the acquired RGB pixel value.
This design allows our acquisition process to be fully differentiable, thus suitable as a forward process in a neural field flow. In this setup, the depth-dependent PSF kernels are pre-defined and we only learn the interpolation weights between the fixed set of produced images. See Figure 1 for visual illustration. The differentiable variant allows us to optimize the whole system in a self-supervised manner, where the generator first maps the input encoding to a pair of intermediate outputs (an all-in-focus image and a pixel-level depth map) that are fed as inputs to the DCM in order to reproduce the acquired image taken by the camera.
The joint optimization of producing both an all-in-focus image and a pixel-level depth map forces an alignment between them, also the pass in the DCM forces the intermediate outputs not to collapse to a trivial blurry solution but to an accurate solution and yields better results compared to previous works where a model was trained to produce just a depth map or just an all-in-focus image.
To optimize our neural network weights to map a coded input to pixel values we used the non-differential acquisition simulation to produce coded-phase images to act as our supervision. Our objective function is to the minimize difference between a given image and the forward process output. Similarly to [13], we first use L2 as our reconstruction loss and switch to SSIM [41] after a fixed number of iterations (see Section 4.2 for ablation on the benefits of the switch). We can formulate our optimization objective as follows:
$\theta^{*}=\operatornamewithlimits{argmin}_{\theta}\ell(h(f_{\theta}(z))-y),$ (2)
4 Experimental Results
4.1 Simulation Results
To evaluate the performance of our method quantitatively both in terms of depth estimation and all-in-focus image reconstruction, we first perform an evaluation on a simulated dataset, namely, ‘TAUAgent’ dataset [^1] https://www.cs.toronto.edu/ harel/TAUAgent/home.html [4]. The Agent dataset consists of a total of 530 images from five synthetic scenes created using the ‘Blender’ computer graphics software (named “City”, “Headbutt”, “Sitting”, “WallSlam”, “WuManchu”). Each scene consists of a ground-truth low-noise all-in-focus image, along with its corresponding pixel-wise accurate depth map. Such data enables an exact depth-dependent imaging simulation, with the corresponding depth of focus (DOF) effects. In contrast to prior works [7, 4, 3], we do not use the ‘ground-truth’ data from the Agent dataset in our optimization. We only use the acquired coded images produced by the camera simulation. The all-in-focus GT images and the GT depth maps are only used to evaluate the performance of our method. For the all-in-focus evaluation, we used PSNR and SSIM, while for the depth estimation, we used average error in meters (as done in [3, 4]). To make a fair comparison we simulate the imaging with the same optical parameters as done in [7, 3, 4].
Note that a crucial disadvantage of previous works is the high dependency on a large and high-quality dataset for successful learning. High-quality datasets like the ‘agent’ dataset are usually small and cumbersome to collect, thus, making the learning much more challenging. In contrast to them, our method reduces this constraint since we explicitly incorporate the DCM into the learning scheme. Thus, we can extract both the all-in-focus RGB image and the pixel-level depth map from a single captured image. We took a subset of 110 images (namely, “agent subset”) from the scenes mentioned above to evaluate the performance on several tasks namely passive depth estimation, image reconstruction, and image deblurring. Lastly, we also show the effect of the implicit generator on the performance of the mentioned tasks. We ran all the experiments on RTX-2090Ti. Each sample took $\sim$ 15 minutes and the memory footprint is 7.5GB for $1024\times 512$ RGB images.
4.1.1 Passive Depth Estimation
To assess our method’s performance in passive depth estimation we compared ourselves to the monocular depth estimation network for phase-coded imaging from [4] which built upon the same phase-coded imaging system. We also followed the same optical parameters to produce the same acquired images. The evaluation was done on the relevant part of the “agent subset” (“WuManchu” scene) that was not used for training [4]. Table I shows a quantitative comparison in terms of RMSE error in meters. Visual results can be found in Figure 3. Overall, our method outperformed the mono network by two orders of magnitude and effectively managed to produce an accurate pixel-level depth map solely from a given captured image. More examples can be found in the supplementary materials.
@@ 4.1.1 Passive Depth Estimation
TABLE I: A quantitative comparison of depth estimation. Results are reported in terms of depth RMSE error [m] ($\downarrow$). Our method outperforms the mono network from [4] in depth estimation on the same imaging system, even though the Monocular depth estimation network from [4] was trained in a supervised manner on the same dataset. This is likely due to the fact that the dataset is small and the model is prone to overfitting.
@@
Supervised 0-shot
GT Input Mono from [4] Error map (Mono-GT) DPCIP Error map (DPCIP-GT)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 3: A qualitative comparison of depth map estimation reveals that our method generates more precise depth maps compared to the mono network from [4], In order to see the performance gap between the methods, we added an absolute difference error map for each method, we can see that DPCIP tends to produce much more accurate depth maps in the background and on the main object compared to the mono network. Note that the depths are clipped to the imaging system’s physical range.
4.1.2 Extended Depth Of Field (EDOF)
To assess our method performance in the task of extended depth of field (EDOF) we compared ourselves to [7] which built upon the same phase-coded imaging system. We also follow the same optical parameters to produce the same acquired images. In this case, the original method was trained on a different synthetic dataset, so we could test all the images in our subset. Table II shows a quantitative comparison in terms of PSNR and SSIM. Visual results can be found in Figure 4. Our method outperformed the baseline by $5dB$ in PSNR and 0.13 in SSIM on average and showed an improvement in terms of image reconstruction metrics. To better show our method’s advantages we did additional effort optimizing EDOF architecture and hyperparameters. We were able to improve from the original work by $\sim 2dB$ on average. Yet, we were still $\sim 2dB$ on average below our zero-shot method. Overall our method outperformed the original EDOF baseline [7] and our improved EDOF baseline both in terms of PSNR [dB] and SSIM. Remarkably, relying only on a single given phase-coded image. Table II shows a quantitative comparison using the agent subset. Qualitative results can be found in Figure 4. More examples can be found in the supplementary materials.
TABLE II: A quantitative comparison of image reconstruction (PSNR [dB] $\uparrow$/ SSIM $\uparrow$). Our method outperforms the existing supervised baseline significantly. The PSNR average improvement of our method is $\sim 5dB$. We re-trained EDOF with an improved architecture and optimization process. DPCIP performs better in $\sim 2dB$ over the improved supervised baseline.
Supervised Supervised 0-shot
Test Scene Original EDOF Improved EDOF DCPIP
City 24.01/0.65 28.75/0.82 29.73/0.91
Headbutt 28.54/0.92 27.673/0.8346 31.28/0.94
Sitting 26.12/0.81 29.12/0.84 30.04/0.91
WallSlam 26.94/0.87 32.54/0.91 31.45/0.93
WuManchu 23.67/0.71 26.546/0.71 30.86/0.93
Supervised 0-shot
GT Input EDOF [7] DPCIP